# Common Trends and Long-Run Multipliers in Nonlinear Structural VARs

### v1

@@ Abstract
While it is widely recognised that linear (structural) VARs may omit important features of economic time series, the use of nonlinear SVARs has to date been almost entirely confined to the modelling of stationary time series, because of a lack of understanding as to how common stochastic trends may be accommodated within nonlinear models. This has unfortunately circumscribed the range of series to which such models can be applied – and/or required that these series be first transformed to stationarity, a potential source of misspecification – and prevented the use of long-run identifying restrictions in these models. To address these problems, we develop a flexible class of additively time-separable nonlinear SVARs, which subsume models with threshold-type endogenous regime switching, both of the piecewise linear and smooth transition varieties. We extend the Granger–Johansen representation theorem to this class of models, obtaining conditions that specialise exactly to the usual ones when the model is linear. We further show that, as a corollary, these models are capable of supporting the same kinds of long-run identifying restrictions as are available in linearly cointegrated SVARs.
@@ 1 Introduction
has played a central role in empirical macroeconomics. Structural VARs provide a satisfyingly coherent framework within which the problem of identifying causal relations, in the presence of simultaneity, may be approached by regarding the $p$ endogenous variables $z_{t}$ as dependent on an underlying sequence of $p$ exogenous structural shocks $\varepsilon_{t}$. Viewed through the lens of these models, the identification problem reduces to one of identifying these structural shocks and their associated impulse responses: equivalently, of identifying the mapping $\Upsilon$ between $\varepsilon_{t}$ and $u_{t}$.
@@
The early literature on structural VARs, following the seminal contribution of Sims (1980), developed contemporaneously with some major advances in the modelling of common (stochastic) trends in macroeconomic time series. This latter work culminated in the development of the theory of the cointegrated VAR, and in particular the Granger–Johansen representation theorem (GJRT), which showed how a VAR could be configured so as to accommodate the presence of common stochastic trends, and the cointegrating relations that are dual to them (Engle and Granger, 1987; Johansen, 1991). This demonstrated that the SVAR model could be reconciled with some of the evident properties of the (levels of) nonstationary macroeconomic time series, obviating the need to difference these series to stationarity prior to analysis (indeed, an implication of this work was that such pre-filtering was not only unnecessary, but could induce model misspecification). An important corollary to the GJRT was that by uncovering the mapping between the common trends and the endogenous variables, it made available a class of long-run identifying restrictions, relating to whether certain structural shocks were regarded as having permanent, or merely transient, effects (Blanchard and Quah, 1989; King et al., 1991).
@@
Some subsequent research on the (S)VAR model has sought to extend it so as to allow for various forms of nonlinearity: typically as modelled via some sort of regime switching, possibly with smooth transitions between these regimes (see e.g. Chan, 2009; Teräsvirta et al., 2010; Hubrich and Teräsvirta, 2013). Recent interest in macroeconomic models that incorporate the possibility of occasionally binding constraints, of which the zero lower bound on interest rates is a leading instance, has its counterpart in the development of new class of regime-switching nonlinear VARs. These newer models are notably distinguished from that earlier literature by being able to accommodate endogenous regime switching, i.e. of allowing the current regime to depend on the current values of the endogenous variables, rather than being pre-determined, or being dependent on some exogenous switching process (Mavroeidis, 2021; Aruoba et al., 2022).
@@
However, there remains a significant disconnect between the literature on nonlinear (S)VARs, and that on cointegration and common trends. In this respect, the situation is reminiscent of that of the literature on linear VARs in the early 1980s, with there being little theoretical understanding of how common trends may be accommodated within nonlinear VARs, such that the removal of common trends by pre-filtering remains a necessity. One area where significant efforts have been made to remedy this pertain to a certain class of ‘nonlinear VECM’ models. These are VAR models that are specified directly in VECM form, with a linear cointegrating space, but in which the stationary components of the model (the equilibrium errors and the lagged differences) are permitted to enter nonlinearly (see e.g. Escribano and Mira, 2002; Bec and Rahbek, 2004; Saikkonen, 2005, 2008; Kristensen and Rahbek, 2010, 2013). This class of models has proved sufficiently tractable to facilitate a nonlinear extension of the GJRT, but the assumptions made are such as to leave a substantial portion of the space of nonlinear VAR models unexplored.
@@
where $f_{0}:\mathbb{R}^{p}\rightarrow\mathbb{R}^{p}$ is continuous and has a continuous inverse (i.e. it is a homeomorphism), and each $f_{i}:\mathbb{R}^{p}\rightarrow\mathbb{R}^{p}$ is continuous. We aim not only to extend the GJRT, but also to preserve exactly the kind of long-run identifying restrictions that are yielded by the GJRT in a linear setting, such that it remains possible to fruitfully discriminate between structural shocks on the basis of the permanence (or transience) of their effects. Through an associated VECM representation, this leads us to consider VARs for which the image of the map
@@
is an $r$-dimensional subspace of $\mathbb{R}^{p}$, a requirement that we term the common row space condition (CRSC); effectively, this entails a decomposition of the form $\pi(z)=\alpha\theta(z)$. This reverses a key assumption of the nonlinear VECM literature, where $\theta(z)=\beta^{\mathsf{T}}z$ is a fixed linear function, but $\alpha$ is allowed to vary (as a function of $\beta^{\mathsf{T}}z_{t-1}$ and $\Delta z_{t-1}$, or their lags). As we discuss, the principal motivation for the CRSC is that it ensures that only $q=p-r$ structural shocks may have permanent effects – whereas if the CRSC is violated, all $p$ structural shocks may have permanent effects, even though $z_{t}$ is driven by only $q$ common stochastic trends.
Our conditions on the nonlinear SVAR (LABEL:SVAR-intro) are otherwise rather general, and unlike the preceding literature allow for common stochastic trends to enter $\{z_{t}\}$ nonlinearly, thereby giving rise to nonlinear cointegrating relations between elements of $\{z_{t}\}$. (A rigorous discussion of ‘cointegration’ in this nonlinear setting is deferred to a companion paper, in which the limiting distribution of the standardised process $n^{-1/2}z_{\lfloor n\lambda\rfloor}$ is derived; for a discussion of this in the context of the censored and kinked SVAR, see Duffy et al., 2023.) These conditions, along with a brief discussion of how they connect with the existing nonlinear VECM literature, are presented in 2. We further show that our conditions reduce to exactly those of the linear cointegrated VAR, when each $\{f_{i}\}_{i=0}^{k}$ is linear: and thus they do not appear to be particularly restrictive.
The principal contribution of this paper is to provide an extension of the GJRT to the setting of a nonlinear VAR (3), demonstrating how the model (LABEL:SVAR-intro) may be configured so as to generate common stochastic trends. This in turn permits an analysis of the stability of the steady state equilibria of the model, and a characterisation of the directions in which the shocks $u_{t}$ have only transient effects (4), and thereby yields identifying restrictions on $\Upsilon$. The GJRT also facilitates the calculation of the implied long-run multipliers associated with the shocks (5). We subsequently return, in 6, to a discussion of the role played by the CRSC in ensuring the availability of long-run identifying restrictions. Finally, 7 elaborates on a class of models in which each $f_{i}$ is piecewise affine (or a smooth, convex combination of piecewise affine functions) showing how the abstract conditions imposed on the model may be specialised for this class of models. 8 concludes. Proofs of all results appear in the appendices.
@@ 2 Model
where $f_{i}:\mathbb{R}^{p}\rightarrow\mathbb{R}^{p}$ for $i\in\{0,1,\ldots,k\}$, and $\{u_{t}\}_{t\in\mathbb{Z}}$ is a sequence of ‘innovations’ taking values in in $\mathbb{R}^{p}$. (As a convenient location normalisation, we shall suppose throughout that $f_{i}(0)=0$ for all $i\in\{0,\ldots,k\}$.) By defining
@@
In macroeconomics, a conventional starting point for empirical work is the structural (linear) VAR model, in which the observed series $\{z_{t}\}$ are assumed to be generated from an underlying i.i.d. sequence of structural shocks $\{\varepsilon_{t}\}$, also of dimension $p$, whose elements are mutually orthogonal (or independent), and each of which have an economic interpretation (as e.g. an aggregate supply shock, a monetary policy shock, etc.). In the present setting, this amounts to positing that that there exists a parametrisation $(\tilde{c},\{\tilde{f}_{i}\}_{i=0}^{p})$ of the model (LABEL:nlVAR) for which
$\tilde{f}_{0}(z_{t})=\tilde{c}+\sum_{i=1}^{k}\tilde{f}_{i}(z_{t-i})+\varepsilon_{t}.$
For the purposes of this paper, we will suppose – via restrictions on the parameter space for $\{f_{i}\}_{i=0}^{p}$ and/or the distribution of $u_{t}$ – that $\varepsilon_{t}$ is identified up to an unknown invertible linear transformation $\Upsilon$ of $u_{t}$, i.e.
$\varepsilon_{t}=\Upsilon^{-1}u_{t}$ (2.5)
where $\Upsilon\Upsilon^{\mathsf{T}}=\operatorname{var}(u_{t})$. (Sufficient conditions for (2.5) to hold will be developed in detail elsewhere; these are not quite so trivial as in the linear case, since we now allow each $f_{i}$ to belong to some class of nonlinear transformations.) By analogy with the linear model, we henceforth term $u_{t}$ the reduced form innovations, even though they do not actually correspond to the innovations of a ‘reduced form model’ in the usual sense of the term. [^1] The principal implication of (LABEL:red2struct) here is that, for the purposes of calculating the long-run effects of the structural shock – the structural long-run multipliers – it suffices to compute the corresponding effects of the reduced form shocks, and then postmultiply these by $\Upsilon^{-1}$. By focusing on the latter, as we do in this paper, we can separate the analysis of the model’s dynamics from the identification of the structural shocks.
2.2 The common row space condition
In the linear VAR, where $\pi(z)=\Pi z$, the key assumption required for the model to generate common stochastic trends (i.e. cointegrated $I(1)$ processes) is that $\operatorname{rk}\Pi=r=p-q$, where $q$ is the number of unit roots in the autoregressive polynomial. Equivalently, this condition may be expressed in terms of the image of the map $z\mapsto\Pi z$,
$\operatorname{im}\Pi=\{\Pi z\mid z\in\mathbb{R}^{p}\}=\operatorname{sp}\Pi$
being an $r$-dimensional (linear) subspace of $\mathbb{R}^{p}$. In extending the GJRT to the setting of the nonlinear VAR (LABEL:nlVAR), the role of $\Pi$ (or more precisely, $\Pi z$) will now be played by $\pi(z)$, and our high-level assumptions below will ensure that
@@ 2.2 The common row space condition
remains an $r$-dimensional (linear) subspace of $\mathbb{R}^{p}$, exactly as in the linear case. We term this the common row space condition (CRSC). (Note that there is no corresponding restriction on $\ker\pi$; in particular, this will not be required to be a $q$-dimensional subspace of $\mathbb{R}^{p}$.)
The existing literature on nonlinear VECM models effectively reverses this condition by requiring that $\ker\pi$ be a $q$-dimensional subspace of $\mathbb{R}^{p}$, without restricting $\operatorname{im}\pi$ (see in particular Kristensen and Rahbek, 2010). Loosely speaking, if we suppose that $\pi$ may be decomposed, for each $z\in\mathbb{R}^{p}$, as
$\pi(z)=\alpha(z)\beta(z)^{\mathsf{T}}$ (2.6)
where $\alpha,\beta:\mathbb{R}^{p}\rightarrow\mathbb{R}^{r}$, then the CRSC effectively entails $\alpha(z)=\alpha$ (if $\beta(z)$ is appropriately normalised), but that $\beta(z)$ is otherwise unrestricted; whereas in the existing literature, $\beta(z)=\beta$ is constant, with the result that there is a globally linear cointegrating space, whereas the models developed here allow the cointegrating relations to be nonlinear. The representation theory in the nonlinear VECM literature is thus strictly complementary to the present work. (Note that even the case where $\beta(z)=\beta$ in our model is not encompassed by that literature, because we allow the dynamics of the implied VECM to depend on the level of $z_{t}$, whereas that literature requires that these be governed entirely by the equilibrium errors $\beta^{\mathsf{T}}z_{t}$ and differences $\Delta z_{t}$.)
The motivation for imposing the CRSC lies in our desire to maintain two key properties of the linear structural VAR. Recall that in that model, common stochastic trends arise because some subset of the structural shocks (say, $q$ in number) have permanent effects on $z_{t}$, while the remaining $r=p-q$ shocks have only transitory effects. This provides not only an underpinning economic explanation for the patterns of long-run co-movement present in the data, but also a fruitful source of long-run identifying restrictions (following Blanchard and Quah, 1989, and King et al., 1991): properties that we would like to preserve as we generalise the cointegrated linear structural VAR in the direction of (LABEL:nlVAR).
However, these properties appear to be highly sensitive to departures from the CRSC, even when the assumption of a globally linear cointegrating space is maintained (i.e. when $\beta(z)=\beta$ for all $z\in\mathbb{R}^{p}$ in (LABEL:factor) above). In models where this condition fails, it will generally be the case that the direction in which an impulse $u_{t}=\delta$ has no permanent effect on $z_{t}$ will vary with the magnitude $\delta$. In 6 below, we provide a simple example to illustrate how this may imply that all structural shocks must have a permanent effect on $z_{t}$, even though $z_{t}$ has only $q$ common stochastic trends. Thus, if we want to work with a class of nonlinear VAR models in which (structural) shocks may be distinguished on the basis of their long-run effects – and in which, as a corollary, common stochastic trends arise because only a subset of these shocks have permanent effects – then it would appear that the CRSC cannot be easily dispensed with.
@@ 2.3 High-level assumptions
To state our main assumptions on the model (LABEL:nlVAR), we first provide a more compact representation for (LABEL:VECM-first); its proof appears in B.
@@
$\begin{bmatrix}\Delta f_{0}(z_{t})\\ \Delta\boldsymbol{\zeta}_{t-1}\end{bmatrix}$ $=\begin{bmatrix}c\\ 0_{p(k-1)}\end{bmatrix}+\begin{bmatrix}\pi(z_{t-1})+g_{1}(z_{t-1})\\ \boldsymbol{g}(z_{t-1})\end{bmatrix}+\begin{bmatrix}0&D_{1}\\ 0&D\end{bmatrix}\begin{bmatrix}z_{t-1}\\ \boldsymbol{\zeta}_{t-2}\end{bmatrix}+\begin{bmatrix}u_{t}\\ 0_{p(k-1)}\end{bmatrix}$ (2.19)
@@
(i). The preceding holds exactly as stated when $k\geq 3$; the model with $k\in\{1,2\}$ may be handled as a special case of $k=3$ with $f_{3}=0$ and (possibly also) $f_{2}=0$. Alternatively, we may specialise the above to $k\in\{1,2\}$ by taking $D$ to be an ‘empty’ matrix when $k=1$, and $D=-I_{p}$ when $k=2$. (In the proofs and statements of all results, to avoid these kinds of complications, we shall implicitly maintain throughout, without loss of generality, that $k\geq 3$.)
@@
where the final equality holds only in the case of a linear VAR. By comparison with $\vec{z}_{t}$, we have defined $\boldsymbol{z}_{t}$ so that the r.h.s. (LABEL:VECM-system) involves a nonlinear transformation of only $z_{t-1}$, and is otherwise linear in $\boldsymbol{\zeta}_{t-2}$. This permits a major simplification of the problem when $k\geq 2$; for this the assumption of additive separability is crucially important.
We next define a class ${\cal M}_{r}$ of nonlinear VAR models (parametrisations of (LABEL:nlVAR)), that will be shown to be consistent with the presence of $q\coloneqq p-r$ common (possibly nonlinear) stochastic trends in $\{z_{t}\}$, and $r$ ‘cointegrating relations’ between those series. For some $\theta:\mathbb{R}^{p}\rightarrow\mathbb{R}^{r}$, define
@@
    where ${\cal B}\subset\mathbb{R}^{p\times[p(k-1)+r]}$ is closed and $\max_{\boldsymbol{\beta}\in{\cal B}}\lVert\boldsymbol{\beta}\rVert\leq\overline{b}$, and
@@
(i). We allow $f_{0}$ to be non-trivial to permit (linear and nonlinear) structural VARs to be accommodated within the present setting. In the linear case, $f_{0}^{-1}(z)=\Phi_{0}^{-1}z$ is trivially continuous; requiring $f_{0}$ to be a homeomorphism extends this to a nonlinear setting. When deriving the properties of the series generated by the models in class ${\cal M}_{r}$, we shall first rewrite the system in terms of $z_{t}^{\ast}\coloneqq f_{0}(z_{t})$, which is itself generated by a model in class ${\cal M}_{r}^{\ast}$ (see A.2 in A).
(ii). It follows from the preceding conditions, and principally from (iii), that $f_{i}$ is continuous for each $i\in\{1,\ldots,k\}$; indeed, these are Lipschitz when $f_{0}=I_{p}$ (see A.4 in A).
@@
As developed in 3 below, if $\{z_{t}\}$ is generated by a model from class ${\cal M}_{r}$, then $\psi(z_{t})$ and $\theta(z_{t})$ respectively provide a representation for the ‘common trends’ and ‘equilibrium errors’ in the system. Since $\chi$ is guaranteed to be a homeomorphism under our assumptions (see A.4 in A), these maps provide an exhaustive description of $z_{t}$. As a further consequence of the invertibility of $\chi$, we have $\operatorname{im}\theta=\mathbb{R}^{r}$ and hence $\operatorname{im}\pi=\operatorname{sp}\alpha$, an $r$-dimensional subspace of $\mathbb{R}^{p}$; our assumptions thus ensure that the model satisfies the common row space condition discussed in 2.2 above.
@@
(iv). The case $r=0$ corresponds to $c=0$ and $\pi(z)=0$ for all $z\in\mathbb{R}^{p}$. The results in this paper will be seen to cover this case, if the appropriate adjustments are made by deleting $\theta$ and $\mu$ from those objects in which they appear, so that e.g. now $\boldsymbol{\theta}=\boldsymbol{g}$ and $\chi=\psi$. (To keep this paper to a manageable length, we do not treat this case explicitly in our proofs.) At the other extreme, the case $r=p$ corresponds to a system with no common trends, and the VAR is stationary. For this reason, we generally restrict the statements of our results to the case where $r\in\{0,\ldots,p-1\}$.
@@
(v). (iii) is our principal assumption on the stability of the weakly dependent components of the system, which comprise the ‘equilibrium errors’ $\xi_{t}\coloneqq\mu+\theta(z_{t})$, and the ‘lagged differences’ $\Delta\zeta_{j,t}=\sum_{i=j}^{k-1}\Delta g_{i}(z_{t-i+j})$ for $j\in\{1,\ldots,k-1\}$. In a linear cointegrated VAR, these reduce to the familiar $\xi_{t}=\mu+\beta^{\mathsf{T}}z_{t}$ and $\Delta\zeta_{j,t}=\sum_{i=j}^{k-1}\Gamma_{j}\Delta z_{t-i+j}$, which are stationary by the GJRT. Though in the present setting, $\xi_{t}$ and $\{\Delta\zeta_{j,t}\}_{j=1}^{k-1}$ will generally not be stationary, (LABEL:jsr) nonetheless ensures that these are of strictly smaller order than $z_{t}$ itself, exactly as in a linearly cointegrated system. As shown in 2.4 below, in a linear VAR (iii) reduces to precisely the familiar conditions on the roots of the associated autoregressive polynomial; whereas the form in which it is stated here is applicable to models which depart so far from linearity that no useful notion of ‘autoregressive roots’ is available.

### v2

# Common Trends and Long-Run Identification in Nonlinear Structural VARs

@@ Abstract
While it is widely recognised that linear (structural) VARs may fail to capture important aspects of economic time series, the use of nonlinear SVARs has to date been almost entirely confined to the modelling of stationary time series, because of a lack of understanding as to how common stochastic trends may be accommodated within nonlinear models. This has unfortunately circumscribed the range of series to which such models can be applied – and/or required that these series be first transformed to stationarity, a potential source of misspecification – and prevented the use of long-run identifying restrictions in these models. To address these problems, we develop a flexible class of additively time-separable nonlinear SVARs, which subsume models with threshold-type endogenous regime switching, both of the piecewise linear and smooth transition varieties. We extend the Granger–Johansen representation theorem to this class of models, obtaining conditions that specialise exactly to the usual ones when the model is linear. We further show that, as a corollary, these models are capable of supporting the same kinds of long-run identifying restrictions as are available in linearly cointegrated SVARs.
First version: April 2024. We thank B. Beare, J. Dolado, A. Escribano, S. Johansen, Y. Lu, B. Nielsen, B. Rossi, and seminar participants at UC3 Madrid, Oxford, Sydney, the 2024 BSE Summer Forum and the 26th Dynamic Econometrics Conference, for their helpful comments on earlier drafts of this paper.
@@ 1 Introduction
has played a central role in empirical macroeconomics. Structural VARs provide a satisfyingly coherent framework within which the problem of identifying causal relations, in the presence of simultaneity, may be approached by regarding the $p$ endogenous variables $z_{t}$ as generated by an underlying sequence of $p$ exogenous structural shocks $\varepsilon_{t}$. Viewed through the lens of these models, the identification problem reduces to one of identifying these structural shocks and their associated impulse responses: equivalently, of identifying the mapping $\Upsilon$ between $\varepsilon_{t}$ and $u_{t}$.
@@
The early literature on structural VARs, following the seminal contribution of Sims (1980), developed contemporaneously with major advances in the modelling of common (stochastic) trends in macroeconomic time series. This latter work culminated in the development of the theory of the cointegrated VAR, and in particular the Granger–Johansen representation theorem (GJRT), which showed how a VAR could be configured so as to accommodate the presence of common stochastic trends, and the cointegrating relations that are dual to them (Engle and Granger, 1987; Johansen, 1991). This importantly demonstrated that the SVAR model could be reconciled with some of the evident properties of the (levels of) nonstationary macroeconomic time series, obviating the need to difference these series to stationarity prior to analysis (indeed, an implication of this work was that such pre-filtering was not only unnecessary, but could induce model misspecification). An important corollary to the GJRT was that by uncovering the mapping between the common trends and the endogenous variables, it made available a class of long-run identifying restrictions, relating to whether certain structural shocks were regarded as having permanent, or merely transient, effects (Blanchard and Quah, 1989; King et al., 1991).
@@
Subsequent work has sought to extend the (S)VAR model so as to allow for various forms of nonlinearity: typically as modelled via some sort of regime switching, possibly with smooth transitions between these regimes (see e.g. Chan, 2009; Teräsvirta et al., 2010; Hubrich and Teräsvirta, 2013). More recently, interest in structural macroeconomic models that incorporate occasionally binding constraints, of which the zero lower bound on interest rates is a leading instance, has had its counterpart in the development of new class of regime-switching nonlinear VARs. These newer models are notably distinguished from the earlier nonlinear VAR literature by being able to accommodate endogenous regime switching, i.e. of allowing the current regime to depend on the current values of the endogenous variables, rather than being pre-determined, or being dependent on some exogenous switching process (Mavroeidis, 2021; Aruoba et al., 2022).
@@
However, there remains a significant disconnect between the literature on nonlinear (S)VARs, and that on cointegration and common trends. In this respect, the situation is reminiscent of that of the literature on linear VARs in the early 1980s, with there being little theoretical understanding of how common trends may be accommodated within nonlinear VARs, such that the prior removal of common trends by pre-filtering remains a necessity. One area where significant efforts have been made to remedy this pertain to a certain class of ‘nonlinear VECM’ models. These are VAR models that are specified directly in VECM form, with a linear cointegrating space, but in which the stationary components of the model (the equilibrium errors and the lagged differences) are permitted to enter nonlinearly (see e.g. Escribano and Mira, 2002; Bec and Rahbek, 2004; Saikkonen, 2005, 2008; Kristensen and Rahbek, 2010, 2013). This class of models has proved sufficiently tractable to facilitate a nonlinear extension of the GJRT, but the assumptions made in this literature are such as to leave a substantial portion of the space of nonlinear VAR models unexplored.
@@
where $f_{0}:\mathbb{R}^{p}\rightarrow\mathbb{R}^{p}$ is continuous and has a continuous inverse (i.e. it is a homeomorphism), and each $f_{i}:\mathbb{R}^{p}\rightarrow\mathbb{R}^{p}$ is continuous. [^1] For a detailed explanation of why the structural shocks enter (LABEL:SVAR-intro) in an apparently similar manner to (LABEL:svar-intro), see 2.4 below. We aim not only to extend the GJRT, but also to preserve exactly the kind of long-run identifying restrictions that are yielded by the GJRT in a linear setting, such that it remains possible to fruitfully discriminate between structural shocks on the basis of the permanence (or transience) of their effects. Via an associated VECM representation, this leads us to consider SVARs in which the image of the map
@@
is an $r$-dimensional subspace of $\mathbb{R}^{p}$, a requirement that we term the common row space condition (CRSC); effectively, this entails a decomposition of the form $\pi(z)=\alpha\theta(z)$. This reverses a key assumption of the nonlinear VECM literature, where $\theta(z)=\beta^{\mathsf{T}}z$ is a fixed linear function, but $\alpha$ is allowed to vary (albeit only as a function of $\beta^{\mathsf{T}}z_{t-1}$ and $\Delta z_{t-1}$, or their lags). As we discuss, the principal motivation for the CRSC is that it ensures that only $q=p-r$ structural shocks may have permanent effects – whereas if the CRSC is violated, all $p$ structural shocks may have permanent effects, even though $z_{t}$ has only $q<p$ common stochastic trends.
Our conditions on the nonlinear SVAR (LABEL:SVAR-intro) are otherwise rather general, and unlike the preceding literature allow for common stochastic trends to enter $z_{t}$ nonlinearly, thereby giving rise to nonlinear cointegrating relations between elements of $z_{t}$. (A rigorous discussion of ‘cointegration’ in this nonlinear setting is deferred to a companion paper, in which the limiting distribution of the standardised process $n^{-1/2}z_{\lfloor n\lambda\rfloor}$ is derived; for a discussion of this in the context of the censored and kinked SVAR, see Duffy et al., 2023.) These conditions are presented in 2, followed by a discussion of the structural shock identification problem that arises in this model. We further show that our conditions reduce to exactly those of the linear cointegrated VAR, when each $\{f_{i}\}_{i=0}^{k}$ is linear: and thus they do not appear to be particularly restrictive.
The principal contribution of this paper is to provide an extension of the GJRT to the setting of a nonlinear VAR (3), demonstrating how the model (LABEL:SVAR-intro) may be configured so as to generate common stochastic trends. This in turn permits an analysis of the stability of the steady state equilibria of the model, and a characterisation of the directions in which the shocks $u_{t}$ have only transient effects (4.1), and thereby yields long-run identifying restrictions on $\Upsilon$ (4.2). The GJRT also facilitates the calculation of the implied long-run multipliers associated with the shocks (4.3). We subsequently return, in 5, to a discussion of the role played by the CRSC in ensuring the availability of long-run identifying restrictions. Finally, 6 elaborates on a class of models in which each $f_{i}$ is piecewise affine (or a smooth, convex combination of piecewise affine functions) showing how the abstract conditions introduced in 2 may be specialised to this class of models. 7 concludes. Proofs of all results appear in the appendices.
@@ 2.1 Framework: a nonlinear VAR($k$)
where $f_{i}:\mathbb{R}^{p}\rightarrow\mathbb{R}^{p}$ for $i\in\{0,1,\ldots,k\}$, and $\{u_{t}\}_{t\in\mathbb{Z}}$ is a sequence of ‘shocks’ taking values in $\mathbb{R}^{p}$. (As a convenient location normalisation, we shall suppose throughout that $f_{i}(0)=0$ for all $i\in\{0,\ldots,k\}$.) By defining
@@
In the linear VAR, where $\pi(z)=\Pi z$ is linear, the key assumption required for the model to generate common stochastic trends (i.e. cointegrated $I(1)$ processes) is that $\operatorname{rk}\Pi=r=p-q$, where $q$ is the number of unit roots in the autoregressive polynomial. Equivalently, this condition may be expressed in terms of the image of the map $z\mapsto\Pi z$,
$\operatorname{im}\Pi=\{\Pi z\mid z\in\mathbb{R}^{p}\}=\operatorname{sp}\Pi,$
being an $r$-dimensional (linear) subspace of $\mathbb{R}^{p}$. In extending the GJRT to the setting of the nonlinear VAR (LABEL:nlVAR), the role of $\Pi$ (or more precisely, of $z\mapsto\Pi z$) will now be played by $\pi(z)$, and our high-level assumptions below will ensure that
@@
remains an $r$-dimensional (linear) subspace of $\mathbb{R}^{p}$, exactly as in the linear case. We term this the common row space condition (CRSC). (Note that there is no corresponding restriction on $\ker\pi$; in particular, this will not be required to be a $q$-dimensional subspace of $\mathbb{R}^{p}$.) For a further discussion of this condition – in particular, of the role that it plays in ensuring that the long-run identifying restrictions familiar from a linear cointegrated VAR are also available in the present setting – see 5 below.
@@ 2.2 High-level assumptions
To state our main assumptions on the model (LABEL:nlVAR), we first provide a more compact representation for (LABEL:VECM-first); its proof appears in C.
@@
$\begin{bmatrix}\Delta f_{0}(z_{t})\\ \Delta\boldsymbol{\zeta}_{t-1}\end{bmatrix}$ $=\begin{bmatrix}c\\ 0_{p(k-1)}\end{bmatrix}+\begin{bmatrix}\pi(z_{t-1})+g_{1}(z_{t-1})\\ \boldsymbol{g}(z_{t-1})\end{bmatrix}+\begin{bmatrix}0&D_{1}\\ 0&D\end{bmatrix}\begin{bmatrix}z_{t-1}\\ \boldsymbol{\zeta}_{t-2}\end{bmatrix}+\begin{bmatrix}u_{t}\\ 0_{p(k-1)}\end{bmatrix}$
@@
(i). The preceding holds exactly as stated when $k\geq 3$; the model with $k\in\{1,2\}$ may be handled as a special case of $k=3$ with at least one of $f_{2}$ and $f_{3}$ being identically zero. Alternatively, we may specialise the above to $k\in\{1,2\}$ by taking $D$ to be an ‘empty’ matrix when $k=1$, and $D=-I_{p}$ when $k=2$. (In the proofs and statements of all results, to avoid these kinds of complications, we shall implicitly maintain throughout, without loss of generality, that $k\geq 3$.)
@@
where the final equality holds only in a linear VAR. By comparison with $\vec{z}_{t}$, we have defined $\boldsymbol{z}_{t}$ so that the r.h.s. of (LABEL:VECM-system) involves a nonlinear transformation of only $z_{t-1}$, and is otherwise linear in $\boldsymbol{\zeta}_{t-2}$. This permits a major simplification of the analysis of the model when $k\geq 2$; for this the assumption of additive separability is crucially important.
We next define a class ${\cal M}_{r}$ of nonlinear VAR models, that will be shown to be consistent with the presence of $q=p-r$ common stochastic trends in $\{z_{t}\}$, and $r$ ‘cointegrating relations’ between those series. For some $\theta:\mathbb{R}^{p}\rightarrow\mathbb{R}^{r}$, define
@@
    where ${\cal B}\subset\mathbb{R}^{p\times[p(k-1)+r]}$ is closed with $\max_{\boldsymbol{\beta}\in{\cal B}}\lVert\boldsymbol{\beta}\rVert\leq\overline{b}$ and
@@
(i). We allow $f_{0}$ to be non-trivial to permit (linear and nonlinear) structural VARs to be accommodated within the present setting: see 2.4 for a further discussion. In the linear case, $f_{0}^{-1}(z)=\Phi_{0}^{-1}z$ is trivially continuous; requiring $f_{0}$ to be a homeomorphism extends this to a nonlinear setting. When deriving the properties of the series generated by the models in class ${\cal M}_{r}$, we shall first rewrite the system in terms of $z_{t}^{\ast}\coloneqq f_{0}(z_{t})$, which is itself generated by a model in class ${\cal M}_{r}^{\ast}$ (see B.2 in B).
(ii). It follows from the preceding conditions, and principally from (iii), that $f_{i}$ is continuous for each $i\in\{1,\ldots,k\}$; indeed, these functions are Lipschitz when $f_{0}=I_{p}$ (see B.4 in B).
@@
As developed in 3 below, if $\{z_{t}\}$ is generated by a model from class ${\cal M}_{r}$, then $\psi(z_{t})$ and $\theta(z_{t})$ respectively provide a representation for the ‘common trends’ and ‘equilibrium errors’ in the system. Since $\chi$ is guaranteed to be a homeomorphism under our assumptions (see B.4 in B), these maps provide an exhaustive description of $z_{t}$. As a further consequence of the invertibility of $\chi$, we have $\operatorname{im}\theta=\mathbb{R}^{r}$ and hence $\operatorname{im}\pi=\operatorname{sp}\alpha$, an $r$-dimensional subspace of $\mathbb{R}^{p}$; models in class ${\cal M}_{r}$ thus satisfy the common row space condition introduced in 2.1 above.
@@
(iv). The case $r=0$ corresponds to $c=0$ and $\pi(z)=0$ for all $z\in\mathbb{R}^{p}$. The results in this paper will be seen to cover this case, if the appropriate adjustments are made by deleting $\theta$ and $\mu$ from those objects in which they appear, so that e.g. now $\boldsymbol{\theta}=\boldsymbol{g}$ and $\chi=\psi$. (To keep this paper to a manageable length, we do not treat this case explicitly in our proofs.) At the other extreme, the case $r=p$ corresponds to a stationary system with no common trends. For this reason, we generally restrict the statements of our results to the case where $r\in\{0,\ldots,p-1\}$.
@@
(v). (iii) is our principal assumption on the stability of the weakly dependent components of the system, which comprise the ‘equilibrium errors’ $\xi_{t}\coloneqq\mu+\theta(z_{t})$, and the ‘lagged differences’ $\Delta\zeta_{j,t}=\sum_{i=j}^{k-1}\Delta g_{i}(z_{t-i+j})$ for $j\in\{1,\ldots,k-1\}$. In a linear cointegrated VAR, these reduce to the familiar $\xi_{t}=\mu+\beta^{\mathsf{T}}z_{t}$ and $\Delta\zeta_{j,t}=\sum_{i=j}^{k-1}\Gamma_{j}\Delta z_{t-i+j}$, which are stationary (up to initialisation) by the GJRT. Though in the present setting, $\xi_{t}$ and $\{\Delta\zeta_{j,t}\}_{j=1}^{k-1}$ will generally not be stationary, (LABEL:jsr) nonetheless ensures that these are of strictly smaller order than $z_{t}$ itself, just as in a linearly cointegrated system. As noted in 2.3 below, in a linear VAR (iii) reduces to precisely the familiar conditions on the roots of the associated autoregressive polynomial; whereas the form in which it is stated here is applicable to models which depart so far from linearity that no useful notion of ‘autoregressive roots’ is available.